# VIP: Versatile Image Outpainting Empowered by Multimodal Large Language Model

### v2

@@ Abstract.
In this paper, we focus on resolving the problem of image outpainting, which aims to extrapolate the surrounding parts given the center contents of an image. Although recent works have achieved promising performance, the lack of versatility and customization hinders their practical applications in broader scenarios. Therefore, this work presents a novel image outpainting framework that is capable of customizing the results according to the requirement of users. First of all, we take advantage of a Multimodal Large Language Model (MLLM) that automatically extracts and organizes the corresponding textual descriptions of the masked and unmasked part of a given image. Accordingly, the obtained text prompts are introduced to endow our model with the capacity to customize the outpainting results. In addition, a special Cross-Attention module, namely Center-Total-Surrounding (CTS), is elaborately designed to enhance further the the interaction between specific space regions of the image and corresponding parts of the text prompts. Note that unlike most existing methods, our approach is very resource-efficient since it is just slightly fine-tuned on the off-the-shelf stable diffusion (SD) model rather than being trained from scratch. Finally, the experimental results on three commonly used datasets, i.e. Scenery, Building, and WikiArt, demonstrate our model significantly surpasses the SoTA methods. Moreover, versatile outpainting results are listed to show its customized ability. Our source code is available at: https://github.com/ucasyjz/VIP.
@@
Image Outpainting, Large Language Model, Prompt Learning
ccs: Computing methodologies Computer vision problems
Figure 1. The visualization of versatile outpainting results generated by our proposed method. Unlike existing methods (left), our method (right) can generate satisfactory outpainting results in both unconditional and conditional paradigms.
1. Introduction
@@
In recent years, with the tremendous progress in image synthesis field (Yi et al., 2023; Pan et al., 2023; Lu et al., 2023; Qiu et al., 2023), image outpainting has become a striking topic. It aims to generate new content based on the original sub-region of a given image. The extrapolated content of the surrounding regions is expected to be consistent with the original content of the center region. Compared to the image inpainting task (Pathak et al., 2016; Elharrouss et al., 2020; Wang et al., 2023; Yang et al., 2017), outpainting is more challenging due to the lack of adequate contextual information between the extrapolated image region and the original center image region.
@@
In the early stage, a line of methods is devoted to improving image outpainting via structure modification of the base model. For example, SRN (Wang et al., 2019) introduces semantic content re-generation based on the Convolutional Neural Network (CNN) architecture. Then, U-Transformer (Gao et al., 2023) is proposed by replacing the CNN structure with the transformer structure to obtain better results. To address the issue of object outlines in extrapolated regions not aligning with the input sub-images, (Gao et al., 2024) exploit eural Ordinary Differential Equations (Neural ODEs) to build a unique bottleneck architecture that enables seamless extrapolation. With its prevalence in numerous visual generation tasks, there have been recent works attempting to introduce diffusion models (Ho et al., 2020) to resolve the image outpainting problem. PQdiff (Zhang et al., 2024b) learns the correlation between different image regions using the diffusion pipeline.
Although previous works are built based on various architectures, the commonality of them is they all assume the given sub-image should be the sole input for image outpainting as query condition. We consider this assumtion severely limits their application and scalability in real scenarios. Firstly, in comparison to the single-variety output, it is more convenient to generate multiple outpainting results according to the preference of user. Second, most existing approaches are trained from scratch tailored for the outpainting task, which have not fully unleashed the potential capacity of the off-the-shelf large generation models, such as SD (Rombach et al., 2022) and DALLE (Ramesh et al., 2021). By contrast, the paradigm of fine-tuning the large generation models for specific tasks has been proven effective in plentiful related works (Kawar et al., 2023; Karras et al., 2023; Everaert et al., 2023; Zhang et al., 2023a; Moon et al., 2022).
To resolve this problem, we present to introduce the text prompts as semantic guidance of the image outpainting pipeline. Doing so allows us to customize the outpainting results in accordance with our willingness. Moreover, the intervention of text prompts facilitates our exploitation of the well-pretrained text-to-image generation model (Rombach et al., 2022; Ramesh et al., 2022; Saharia et al., 2022).
@@
Besides, in existing works (Zhang et al., 2024b; Yao et al., 2022) for image outpainting, the relation modeling between different image blocks is conducted in random regions. Actually, we argue there is a nonnegligible distinction between the central and surrounding regions of an image. Specifically, in the center regions, the content tends to be noteworthy objects or primary information, while in surrounding regions, the content tends to be background information. Considering image outpainting extends from the central regions to the surrounding regions, it is imperative to capture the semantic association between these two spatial positions explicitly.
@@
This paper addresses the above issues by proposing a novel framework namely Versatile Image Outpainting (VIP). First, to enable this model to be aware of the space information through text prompt as query instruction, we need to construct the pairwise prompt-image data. To achieve it, we leverage the Multimodal Large Language model (MLLM) to produce the textual descriptions of the central and surrounding regions of images, followed by organizing the results into a unified format for output. These operations were performed automatically without the need for manual annotation. Based on the obtained image-prompt data pair, to better leverage the capabilities of the pre-trained large model, we construct our image outpainting model modified from a off-the-shelf stable diffusion (SD) model (Rombach et al., 2022). Furthermore, we elaborately design a Center-Total-Surrounding (CTS) cross-attention module to capture the fine-grained correspondences between the central and surrounding regions of the image and the corresponding parts in the text prompt, which enhances the customization and generalization capabilities of our model. Notably, unlike existing methods, our model is computationally efficient, requiring only fine-tuning on a small amount of data to achieve the current best outpainting effects. As illustrated in Fig. 1, we show that the introduction of the novel text prompt format not only helps us achieve diverse customization results, but also significantly improves the quality of the extrapolated images.
@@
  • We build a novel diffusion model based image outpainting framework that is characterized by customizing the results according to the requirements of users. By leveraging the Multimodal Large Language model to annotate images with a novel prompt format automatically, we construct our model via fine-tuning the pre-trained stable diffusion model.
@@
  • A novel Center-Total-Surrounding (CTS) cross-attention module is presented to promote the fine-grained interaction between corresponding parts in images and text prompts.
@@
  • The experimental results not only validate our approach substantially outperforms the existing methods, but also can generate diverse results based on customized instructions.
2. Related Work
Figure 2. The overall framework of our proposed learning method. First, we concatenate the raw image latent, masked image latent, and mask as the input of the unet. The output will be calculated as mean-square error (MSE) loss with the added noise in the raw image latent. Meanwhile, the proposed text prompt and mask will be input into the proposed cross-attention module as conditions. ’R’ and ’C’ mean the resize and concatenation operations, respectively.
2.1. Image Outpainting
@@
The task aims to extrapolate the surrounding contents based on the center contents of the image. The research (Sabini and Rusak, 2018) introduces the deep learning method into the image outpainting task but only performs the horizontal outpainting. IOH (Van Hoorick, 2019) is proposed by utilizing the Generative Adversarial Network (GAN) (Creswell et al., 2018) to implement the image outpainting. SRN (Wang et al., 2019) is developed by directly learning the semantic features for outpainting. Kim et al. (Kim et al., 2021) transform the outpainting task into the inpainting task by swapping the outer region of the image with its inner region. Lu et al. (Lu et al., 2021) propose a bidirectional content transfer module to generate the intermediate region between two different input images. Gao et al. (Gao et al., 2023) further improve the results with a U-transformer structure and extend the horizontal outpainting to all-side outpainting. QueryOTR (Yao et al., 2022) further adopts the ViT module and transforms this task into a patch-wise sequence-to-sequence autoregression problem. Neural ODE (Gao et al., 2024) is designed by implementing continuous extrapolation in latent space to make the generation results smoother. However, these methods are all trained without text prompt conditions, which suffer from the lack of semantic information guidance. Note that Inout (Cheng et al., 2022) also introduces text as conditions to control the outpainting results. Its text prompt annotation is just represented as one label word. This label is annotated by a semantic segmentation model that only performs close-vocabulary classification on a narrow range of classes. By contrast, our employed MLLM could return the open-world results and automatically organize them into templated answers. Moreover, Inout is built based on GAN architecture rather than diffusion models as ours, which limits its capacity.
2.2. Diffusion Model
The diffusion model (Dhariwal and Nichol, 2021; Avrahami et al., 2022; Zhang et al., 2024a; Bar-Tal et al., 2024; Yang et al., 2023a; Zhou and Tulsiani, 2023) is a generative model for producing high-quality, diverse data, such as images, video, and text. These models generate data by gradually introducing noise into the data and then learning how to reverse the process.
DDPM (Ho et al., 2020) lays the foundation of the diffusion era for the generation task. It generates data by simulating the diffusion process of data with high quality and diversity. To speed up the diffusion process by reducing computation costs, LDM (Rombach et al., 2022) transforms the image space into the latent space through a vae encoder and maps it back to the image space using a vae decoder. Meanwhile, it incorporates the text condition into the pipeline for customized generation. Based on LDM, Stable Diffusion is a commonly-used pretrained large-scale model, which has extensively promoted the development of the generation field. ControlNet (Zhang et al., 2023b) is proposed to add conditional control by copying a trainable version based on stable diffusion. DALLE-2 (Ramesh et al., 2022) combines the clip and diffusion model to obtain better text-to-image results. Dreambooth (Ruiz et al., 2023) achieves subject-driven generation by fine-tuning the pretrained stable diffusion model on just a few images of a subject. PQDiff (Zhang et al., 2024b) is proposed by learning the correlation between randomly cropped views for image outpainting based on a diffusion model. However, it still lacks text conditions.
2.3. Multimodal Large Language model
@@
Multimodal Large Language models are designed to process and understand multiple types of data (such as text, images, video, etc.). These models can learn across different data modes and capture associations between different modes to perform well on a range of cross-modal tasks, including Visual Question Answering (VQA), multimodal translation, and more. Here are some important multimodal Multimodal Large Language models. GPT-4 (Yang et al., 2023b) is an open-source MLLM that can understand the image. With the Reinforcement Learning from Human Feedback (RLHF), the feedback from humans enhances the output results. Llama2 (Touvron et al., 2023) is proposed by FaceBook with some improvements in structure and data training. Through three-stage training, Qwen-VL (Bai et al., 2023) further enhances the visual ability of MLLM by introducing the visual encoder. ERNIE-ViL (Yu et al., 2021) adopts scene graphs of visual scenes to build detailed semantic connections across vision and language.
3. Methods
@@
3.1. Preliminaries
Diffusion Model Diffusion Models are a class of deep learning models that generate high-quality data by gradually increasing and then decreasing the noise on the data. This model performs well in the fields of image generation, video generation, text generation, etc. The working principle of the diffusion model can be divided into two main stages: the forward and the reverse process. The forward process is demonstrated as follows:
(1) $q_{t}(x_{t}|x_{t-1})=N(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$
@@ 3.1. Preliminaries
(2) $x_{t-1}=\frac{1}{1-\beta_{t}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{a}_{t}}}\epsilon_{\theta}(x_{t},t))$
@@
(3) $\mathcal{L}_{DM}=\mathbb{E}_{x_{0},\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\right]$
Multimodal Large Language model Multimodal Large Language models (MLLM) are advanced artificial intelligence technique that combines the ability to process different types of data. These models gain the ability to understand and generate content across multiple modals by learning large multimodal datasets. They can understand the relationship between text and images, generate text prompts from images, etc. The users should give an input prompt to the MLLM for usage. For example, ’description of the input image’, and the MLLM will become an expert in the image caption field and generate the corresponding description. In this paper, we utilize the GPT-4 (Yang et al., 2023b) model due to its excellent performance.
Figure 3. The architecture of Center-Total-Surrounding Cross-attention module. $\cdot$’, ’$\times$’, and ’+’ mean the element-wise multiplication, tensor multiplication and element-wise addition, respectively.
3.2. Data Generation
@@
Specifically, we define the task prompt as follows, [Please use three English keywords to describe the center and the surrounding of the image, and output in the following format:‘Center:xxx,xxx,xxx; Surrounding:xxx,xxx,xxx’]. The MLLM takes this prompt as input and auto-regressively captions the different regions of the input image. Then, the data pairs of the enhanced text prompt with space information and the corresponding image are obtained. After generation, we filter the data to reduce the noisy samples and set a part of the data with a prompt ’Center:;Surrounding’ without keywords as the unconditional outpainting, and the other parts with keywords as the condition outpainting. Therefore, customized outpainting can be achieved based on the generation of novel text conditions. Then, the corresponding fine-tuning can be performed using the proposed text prompt format.
3.3. Center-Total-Surrounding Cross-Attention
In this section, we will describe the implementation of the Center-Total-Surrounding (CTS) Cross-Attention module using the proposed text prompt format. To achieve the alignment between different image regions and the corresponding text parts, the text prompt part for a specific image area should be input into the model, independently. Therefore, we propose a CTS Cross-Attention to realize the purpose as shown in Fig. 3. Specifically, the details of the implementation are described below.
@@ 3.3. Center-Total-Surrounding Cross-Attention
(4) $Q$ $=W_{Q}\cdot F_{I},K=W_{K}\cdot F_{T},V=W_{V}\cdot F_{T}$
@@
However, it can be found that the space information is lacking in the original text prompt, and the different text tokens cannot correspond accurately to different image tokens in the masked image without any guidance. Therefore, the CTS cross-attention is proposed to generate the results of the local cross-attention for better conditions. Specifically, we first split the prompt into two parts given a prompt as ’Center:xxx; Surrounding:xxx’ according to the center and surrounding settings. Then, we develop two additional cross-attention modules to control the generation of the center and surrounding regions.
@@
(5) $Y^{C}$ $=\text{softmax}\left(\frac{{Q{K^{C}}^{T}}}{{\sqrt{d_{k}}}}\right)V^{C}$
@@
Table 1. Comparison results with the state-of-the-art methods. The best and second best results are indicated by boldface and underline, respectively. ’+ copy’ means that after the network generates extrapolated images, the model duplicates the input sub-images into the corresponding central regions of the extrapolated images.
@@
SRN (Wang et al., 2019) GAN-based 47.781 2.981 38.644 3.862 76.749 3.629
NSIPO (Yang et al., 2019) GAN-based 25.977 3.059 30.465 4.153 22.242 5.600
IOH (Van Hoorick, 2019) GAN-based 32.107 2.886 49.481 3.924 40.184 4.835
U-Transformer (Gao et al., 2023) GAN-based 20.575 3.249 30.542 4.189 15.904 6.567
Neural ODE (Gao et al., 2024)+copy GAN-based 19.352 3.998 21.009 5.317 14.828 8.957
QueryOTR(Yao et al., 2022)+copy GAN+MAE 20.366 3.955 22.378 4.978 14.955 7.896
PQdiff (Zhang et al., 2024b) Diffusion-based 29.446 3.849 28.855 4.879 10.545 7.374
PQdiff (Zhang et al., 2024b) +copy Diffusion-based 20.100 3.981 19.133 5.350 7.968 8.605
@@
The initial weights of the two new attention modules are copied from the original attention module to maintain the strong generation ability of the pretrained model. Meanwhile, based on the input mask information, we can separate the latent features of the center and surrounding regions. Then, the final results are obtained by adding the two parts:
(6) $\hat{Y}^{C}$ $=Y^{C}*(1-mask)$
$\hat{Y}^{S}$ $=Y^{S}*mask$
$\hat{Y}$ $=\hat{Y}^{C}+\hat{Y}^{S}$
(7) $Y$ $=(1-a)*Y+a*\hat{Y}$
4. Experiment
4.1. Dataset

### v3

@@ Abstract
In this paper, we focus on resolving the problem of image outpainting, which aims to extrapolate the surrounding parts given the center contents of an image. Although recent works have achieved promising performance, the lack of versatility and customization hinders their practical applications in broader scenarios. Therefore, this work presents a novel image outpainting framework that is capable of customizing the results according to the requirements of users. First of all, we take advantage of a Multimodal Large Language Model (MLLM) that automatically extracts and organizes the corresponding textual descriptions of the masked and unmasked part of a given image. Accordingly, the obtained text prompts are introduced to endow our model with the capacity to customize the outpainting results. In addition, a special Center-Total-Surrounding (C-T-S) decoupled control mechanism is elaborately designed to boost text-driven generation by enhancing the interaction between specific spatial regions of the image and corresponding parts of the text prompts. Note that unlike most existing methods, our approach is very resource-efficient since it is just slightly fine-tuned on the off-the-shelf stable diffusion (SD) model rather than being trained from scratch. Finally, the experimental results on three commonly used datasets, i.e. Scenery, Building, and WikiArt, demonstrate our model significantly surpasses the SoTA methods. Moreover, versatile outpainting results are listed to show its customized ability. Our source code is available at: https://github.com/ucasyjz/VIP
Keywords:
@@
Image Outpainting Multimodal Large Language Model Prompt Learning.
1 Introduction
@@
With significant advancements [47, 22, 20, 24] in image generation, image outpainting has become a striking topic in recent years. This task extends the content of a given image, requiring generated regions to seamlessly align with original regions. Unlike inpainting [23, 7, 40, 43], outpainting is more challenging due to limited contextual information between extrapolated and original image regions.
Figure 1: The visualization of versatile outpainting results generated by our proposed method. Unlike existing methods (left), our method (right) can generate satisfactory outpainting results in both unconditional and conditional paradigms.
@@
In the early stage, a line of methods is devoted to improving image outpainting via structure modification of the base model. For example, SRN [41] introduces semantic content re-generation based on the Convolutional Neural Network (CNN) architecture. Then, U-Transformer [10] is proposed by replacing the CNN structure with the transformer structure for long-range relationship acquisition. To address the issue of object outlines in extrapolated regions not aligning with the input sub-images, Gao et al. [11] exploit eural Ordinary Differential Equations (Neural ODEs) to build a unique bottleneck architecture that enables seamless extrapolation. With its prevalence in numerous visual generation tasks, there have been recent works attempting to introduce diffusion models [13] to resolve the image outpainting problem. PQDiff [51] learns the correlation between different image regions using the diffusion pipeline.
Although previous works are built based on various architectures and obtain satisfactory results, they typically assume that a given sub-image serves as the sole input for image outpainting, limiting their application and scalability in real scenarios. Generating multiple outpainting results based on various user preferences is more practical than producing a single output automatically. Moreover, many existing approaches are trained from scratch for specific scenes, failing to leverage the full potential of pre-trained large generative models like SD [28] and DALLE [26]. In contrast, fine-tuning these large models for specific tasks has been shown to be effective in various studies [15, 14, 9, 52, 21].
@@
To resolve this problem, we present to introduce text prompts as semantic guidance of image outpainting. This approach allows us to customize outpainting results in accordance with our willingness. Moreover, the intervention of text prompts facilitates our exploitation of the well-pre-trained text-to-image generation model [28, 25, 31]. Besides, in existing works [51, 46] for image outpainting, the relation modeling between different image blocks is conducted in random regions. Actually, we argue there is a nonnegligible distinction between central and surrounding regions of an image. Specifically, in center regions, the content tends to be noteworthy objects or primary information, while in surrounding regions, the content tends to be background information. Considering image outpainting extends from the central regions to the surrounding regions, it is imperative to capture the semantic association between these two spatial positions explicitly.
@@
This paper addresses the above issues by proposing a novel framework namely Versatile Image Outpainting (VIP). First, to enable this model to be aware of the space information through text prompt as query instruction, we need to construct the pairwise prompt-image data. To achieve this, we leverage the Multimodal Large Language model (MLLM) to produce the textual descriptions of the central and surrounding regions of images, followed by organizing the results into a unified format for output. These operations are performed automatically without the need for manual annotation. Based on the obtained image-prompt data pair, to better leverage the capabilities of the pre-trained large model, we construct our image outpainting model modified from an off-the-shelf stable diffusion (SD) model [28]. Furthermore, we elaborately design a Center-Total-Surrounding (C-T-S) decoupled control mechanism to capture the fine-grained correspondences between the central and surrounding regions of the image and the corresponding parts in the text prompt, which enhances the customization and generalization capabilities of our model. Notably, unlike existing methods, our model is trained efficient, requiring only fine-tuning on a small amount of data to achieve the current best outpainting effects. As illustrated in Fig. 1, we show that the proposed method not only helps us achieve diverse customization results but also significantly improves the quality of the extrapolated images.
@@
  • Using a pre-trained prior, we develop a novel diffusion-based image outpainting framework for customized results. Images are annotated with an innovative text prompt format automatically by leveraging MLLM.
@@
  • A novel Center-Total-Surrounding (C-T-S) decoupled control mechanism is presented to promote the fine-grained interaction between corresponding parts in images and text prompts.
@@
  • The experimental results not only validate our approach substantially outperforms the existing methods but also can generate diverse results based on customized instructions.
2 Related Work
2.1 Image Outpainting
@@
The task aims to extrapolate the surrounding contents based on the center contents of the image. The research [30] introduces the deep learning method into the image outpainting task but only performs the horizontal outpainting. IOH [39] is proposed by utilizing the Generative Adversarial Network (GAN) [5] to implement the image outpainting. SRN [41] is developed by directly learning the semantic features for outpainting. Kim et al. [16] transform the outpainting task into the inpainting task by swapping the outer region of the image with its inner region. Lu et al. [19] propose a bidirectional content transfer module to generate the intermediate region between two different input images. Gao et al. [10] further improve the results with a U-transformer structure and extend the horizontal outpainting to all-side outpainting. QueryOTR [46] further adopts the ViT module and transforms this task into a patch-wise sequence-to-sequence autoregression problem. Neural ODE [11] is designed by implementing continuous extrapolation in latent space to make the generation results smoother. However, these methods are all trained without text prompt conditions, which suffer from the lack of semantic information guidance. Note that Inout [4] also introduces text as conditions to control the outpainting results. Its text prompt annotation is just represented as one label word. This label is annotated by a semantic segmentation model that only performs close-vocabulary classification on a narrow range of classes. By contrast, our employed MLLM could return the open-world results and automatically organize them into templated answers. Moreover, Inout is built based on GAN architecture rather than diffusion models like ours, which limits its capacity.
2.2 Multimodal Large Language Model
@@
Multimodal Large Language models are designed to process and understand multiple types of data (such as text, images, video, etc.). These models can learn across different data modes and capture associations between different modes to perform well on a range of cross-modal tasks, including Visual Question Answering (VQA), multimodal translation, and more. Here are some important multimodal Multimodal Large Language models. GPT-4 [44] is an open-source MLLM that can understand the image. With the Reinforcement Learning from Human Feedback (RLHF), the feedback from humans enhances the output results. Llama2 [38] is proposed by FaceBook with some improvements in structure and data training. Through three-stage training, Qwen-VL [2] further enhances the visual ability of MLLM by introducing the visual encoder. ERNIE-ViL [48] adopts scene graphs of visual scenes to build detailed semantic connections across vision and language.
2.3 Diffusion Model
The diffusion model [6, 1, 50, 3, 42, 53] is a generative model for producing high-quality, diverse data, such as images, video, and text. These models generate data by gradually introducing noise into the data and then learning how to reverse the process.
DDPM [13] lays the foundation of the diffusion era for the generation task. It generates data by simulating the diffusion process of data with high quality and diversity. To speed up the diffusion process by reducing computation costs, LDM [28] transforms the image space into the latent space through a vae encoder and maps it back to the image space using a vae decoder. Meanwhile, it incorporates the text condition into the pipeline for customized generation. Based on LDM, Stable Diffusion is a commonly-used pre-trained large-scale model, which has extensively promoted the development of the generation field. ControlNet [49] is proposed to add conditional control by copying a trainable version based on stable diffusion. DALLE-2 [25] combines the clip and diffusion model to obtain better text-to-image results. Dreambooth [29] achieves subject-driven generation by fine-tuning the pre-trained stable diffusion model on just a few images of a subject. PQDiff [51] is proposed by learning the correlation between randomly cropped views for image outpainting based on a diffusion model. However, it still lacks text conditions.
Figure 2: Overall framework of our proposed learning method. First, we concatenate the raw image latent, masked image latent, and mask as the input of the unet. The output will be calculated as mean-square error (MSE) loss with the added noise in the raw image latent. Meanwhile, the text prompt and mask will be input into the proposed C-T-S decoupled control mechanism as conditions. ’R’ and ’C’ mean the resize and concatenation operations, respectively.
3 Methods
@@
3.1 Preliminaries
Diffusion Model. Diffusion Models are a class of deep learning models that generate high-quality data by gradually increasing and then decreasing the noise on the data. This model performs well in the fields of image generation, video generation, text generation, etc. The working principle of the diffusion model can be divided into two main stages: the forward and the reverse process. The forward process is demonstrated as follows:
$q_{t}(x_{t}|x_{t-1})=N(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ (1)
@@ 3.1 Preliminaries
$x_{t-1}=\frac{1}{1-\beta_{t}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{a}_{t}}}\epsilon_{\theta}(x_{t},t))$ (2)
@@
$\mathcal{L}_{DM}=\mathbb{E}_{x_{0},\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\right]$ (3)
Multimodal Large Language Model. Multimodal Large Language models (MLLM) are advanced artificial intelligence technique that combines the ability to process different types of data. These models gain the ability to understand and generate content across multiple modals by learning large multimodal datasets. They can understand the relationship between text and images, generate text prompts from images, etc. The users should give an input prompt to the MLLM for usage. For example, ’description of the input image’, and the MLLM will become an expert in the image caption field and generate the corresponding description. In this paper, we utilize the GPT-4 [44] model due to its excellent performance.
3.2 Text Paradigm Incorporated with Space Information
@@
Specifically, we define the task prompt as follows, “Please use three English keywords to describe the center and the surrounding of the image, and output in the following format: Center:xxx,xxx,xxx; Surrounding:xxx,xxx,xxx”. The MLLM takes this prompt as input and auto-regressively captions the different regions of the input image. Then, the data pairs of the enhanced text prompt with space information and the corresponding image are obtained. After generation, we filter the data to reduce the noisy samples and set a part of the data with a prompt ’Center:; Surrounding’ without keywords as the unconditional outpainting, and the other parts with keywords as the condition outpainting. Therefore, customized outpainting can be achieved based on the generation of novel text conditions. Then, the corresponding fine-tuning can be performed using the proposed text prompt format.
3.3 Text-driven Center-Total-Surrounding Decoupled Control
Figure 3: Architecture of Center-Total-Surrounding decoupled control mechanism. $\cdot$’, ’$\times$’, and ’+’ mean the element-wise multiplication, tensor multiplication and element-wise addition, respectively.
In this section, to make full use of the proposed text paradigm and achieve the alignment between different image regions and the corresponding text parts, the text prompt part for a specific image area should be input into the model, independently. It is expected that different spatial positions of images can be generated by corresponding parts of text control. Therefore, we propose a Center-Total-Surrounding (C-T-S) decoupled control mechanism to realize the purpose as shown in Fig. 3. Specifically, the details of the implementation are described below.
@@ 3.3 Text-driven Center-Total-Surrounding Decoupled Control
$Q$ $=W_{Q}\cdot F_{I},K=W_{K}\cdot F_{T},V=W_{V}\cdot F_{T}$ (4)
@@
However, it can be found that the space information is lacking in the original text prompt, and the different text tokens cannot correspond accurately to different image tokens in the masked image without any guidance. Therefore, the C-T-S decoupled control mechanism is proposed to generate the results of the local cross-attention for better conditions. Specifically, we first split the prompt into two parts given a prompt as ’Center:xxx; Surrounding:xxx’ according to the center and surrounding settings. Then, we develop two additional cross-attention modules to control the generation of the center and surrounding regions.
@@
$Y^{C}$ $=\text{softmax}\left(\frac{{Q{K^{C}}^{T}}}{{\sqrt{d_{k}}}}\right)V^{C}$ (5)
@@
The initial weights of the two new attention modules are copied from the original attention module to maintain the strong generation ability of the pre-trained model. Meanwhile, based on the input mask information, we can separate the latent features of the center and surrounding regions. Then, the final results are obtained by adding the two parts:
$\begin{split}\hat{Y}^{C}&=Y^{C}*(1-mask)\\ \hat{Y}^{S}&=Y^{S}*mask\\ \hat{Y}&=\hat{Y}^{C}+\hat{Y}^{S}\end{split}$ (6)
$Y$ $=(1-a)*Y+a*\hat{Y}$ (7)
Table 1: Comparison with State-of-The-Art methods. The best and second best results are indicated by boldface and underline, respectively. ’+ copy’ means that after the network generates outpainting images, input sub-images are duplicated into corresponding central regions of outpainting images.
@@
SRN [41] GAN-based 47.781 2.981 38.644 3.862 76.749 3.629
NSIPO [45] GAN-based 25.977 3.059 30.465 4.153 22.242 5.600
IOH [39] GAN-based 32.107 2.886 49.481 3.924 40.184 4.835
U-Transformer [10] GAN-based 20.575 3.249 30.542 4.189 15.904 6.567
Neural ODE [11]+copy GAN-based 19.352 3.998 21.009 5.317 14.828 8.957
QueryOTR[46]+copy GAN+MAE 20.366 3.955 22.378 4.978 14.955 7.896
PQDiff [51] Diffusion-based 29.446 3.849 28.855 4.879 10.545 7.374
PQDiff [51] +copy Diffusion-based 20.100 3.981 19.133 5.350 7.968 8.605
@@
4 Experiment
4.1 Dataset